# Road Graph Generator: Mapping roads at construction sites from GPS data

### v3

@@ Abstract
We propose a new method for inferring roads from GPS trajectories to map construction sites. This task presents a unique challenge due to the erratic and non-standard movement patterns of construction machinery, which significantly diverge from typical vehicular traffic on established roads. Our proposed method first identifies intersections in the road network that serve as critical decision points, and then connects them with edges to produce a graph, which can subsequently be used for planning and task allocation. We demonstrate the approach by mapping roads at a real-life construction site in Norway. The method is validated on four increasingly complex segments of the map. In our tests, the method achieved perfect accuracy in detecting intersections and inferring roads in data with no or low noise, while its performance was reduced in areas with significant noise and consistently missing GPS updates.
@@ 3 Problem definition
The objective is to use GPS data to infer a graph representing the underlying road network, where each node of the graph represents a road intersection or an action point (loading or off-loading) which is described as a vector of latitude, longitude and altitude $I=(lat,lon,altitude)$. An intersection is defined as a point where at least three roads meet. The edges of the graph represent the roads and are defined as a series of consecutive, equidistant points $E=(lat,lon,altitude)$. An example of a graph is shown in Fig. 3.
@@ 5 Data preprocessing
    The Haversine formula is applied to convert latitude and longitude into coordinates expressed in meters $(x,y)$, with the origin arbitrarily set to the minimum values of latitude and longitude in the data set.
@@
    Each trip is interpolated with a predefined resolution $\Delta D_{\mathrm{inter}}$.
The hyperparameters and their default values are given in the Appendix.
Figure 1: Identification of candidate intersections. Upper plot: The GPS trajectories are converted into movement directions $\Phi\in[0,\pi]$ (bi-directional movement, Sec. 6.1). Bottom plot: The area is split into grid cells, with the median direction determined for each cell. The median direction is then used to compute the directional dissimilarity $\Delta\phi_{i,j}$ between adjacent cells. The cells with $\Delta\phi_{i,j}\geq\Delta\phi_{\text{thr}}$ are identified as intersection candidates.
@@ 6 Graph inference algorithm
In the first step, the map is partitioned into a grid (Fig. 1). Each cell is of size $N_{\mathrm{res}}\times N_{\mathrm{res}}=5\times 5$ m2, and is denoted as $c_{i,j}$, where $i$ and $j$ are the row and column indices, respectively. The GPS trajectories, divided into trips, can then be converted into movement directions $\Phi=\{\phi_{1},\phi_{2},...,\phi_{N}\}$ based on the latitude and longitude of consecutive GPS updates (similar to ijgi4042446). These directions are normalized to a $[0,\pi]$ range (00 to $180^{\circ}$), accounting for bi-directional vehicle movement and treating travel in either direction along the same path as equivalent. Finally, for each set $\Phi_{i,j}$ of points recorded within each cell $c_{i,j}$, the median $\tilde{\phi}_{i,j}$ is computed. This procedure essentially transforms the grid into a 2D histogram, where each cell represents the median of bi-directional vehicular movement.
@@ 6.3 Validation of intersections
The next step consists of filtering out false positives among the candidate intersections using an adapted version of the method proposed in 10.1145/3234692. This method discerns intersections from e.g., road bends by determining the number of roads leaving the candidate location for the intersection. In our version, we first subset the GPS points to those that lie within the circle bounded by the outer radius $R+L$ and apply DBSCAN clustering. The analysis is restricted to the GPS points that are directly connected to the intersection, i.e., they form a cluster with the points passing within a distance of $D_{\mathrm{passing}}=15$ m. Next, following 10.1145/3234692, we define an extremity annulus bounded by two concentric circles with radii $R=30$ m and $R+L=55$ m around each candidate intersection (Fig. 2). These GPS points located within the annulus and are connected to the intersection are clustered with a clustering distance of $D_{\mathrm{ext\_clust}}=30$ m. Each cluster of a size larger than $N_{\mathrm{ext\_clust}}=2$ represents a road that passes through the extremity. The final set of intersections are those that have at least three such clusters in their extremity, representing three roads leaving the intersection.
Depending on the actual size of intersections, this step can be repeated with varying values of $R$ and $L$. To reduce computations for determining distances between the GPS points and intersections, we employed k-dimensional trees (k-d trees) that partition space into a hierarchical tree structure, significantly accelerating spatial queries. This approach allowed us to narrow down calculations to only those pairs of GPS points and intersection centers that are within a predefined maximum distance $R+L=55$ m from each other.
@@ 6.5 Road inference
In the final step of constructing the road network, the nodes, consisting of intersections, loading and drop-off locations, are interconnected by edges that represent the roads. Additional edges are also created to depict the roads with a dead end, i.e., the roads that pass through only one intersection.
@@ 7 Results
Figure 3: Examples of the inferred graph with three types of nodes representing intersections (red), load points (green), and drop-off points (black), and edges representing the roads (blue). The actual GPS data are represented by gray lines.
The visual representation of the algorithm’s performance is presented in Fig. 3. Each of the four subplots shows segments of the inferred graph in the order from relatively straightforward to more complex scenarios. The parameters used are defined in the Appendix. The intersection validation step was run twice, with $R=30$ m and $R=100$ m (Fig. 4).
In the first two examples: Zone 1 and 2, the algorithm achieves perfect recall and precision: All intersections are detected correctly and all driven roads are outlined. The use of at least two radii in the intersection validation step is crucial to detect cross-roads of varying shape and spread, e.g., merging roads (34, Zone 1) and small detours (14, Zone 2). We note that the algorithm is robust against noisy GPS updates present in Zone 1 (visible as straight lines). This robustness against noise that visually resembles a road split is gained when quantifying the directional dissimilarity (Sec. 6.2), as the area is never qualified as an intersection candidate (see upper left in Fig. 4).
In our further analysis, we show the limitation of the algorithm in the presence of consistent noise. In Zone 3 four intersections are detected correctly (7, 8, 9, and 32). Number 33 is falsely identified as an intersection due to the noisy data, while an actual merging of the roads is located 100 m to the left. In this case, the noisy updates are more frequent than the correct outline of the road. Zone 4 presents the biggest challenge to the algorithm, with noisy data dominating several parts of the map, in the areas where the vehicles often change direction. In this case, the algorithm detects several false positive intersections, e.g., number 23, or 30.
@@ 8 Discussion
    Contextual parameter tuning: The algorithm’s parameters have to be tuned for the specific application and characteristics of the landscape. For instance, the distance parameters that define the relevant area surrounding a singular point marked as an intersection will vary depending on the area covered by an intersection: a well-defined crossroad will need tighter parameter settings than more expansive roundabouts or open-area nodes with several diverging routes.
@@
    Quality of the GPS data: While multiple problems with the data can be addressed with the steps outlined in Sec. 5, mitigating certain types of noise is not trivial, especially if the inaccurate GPS updates are dominant and overshadow valid data. Customizing the data cleansing procedure may become essential in specific deployment contexts.
@@ Appendix A Nomenclature
Appendix B Algorithm and data processing parameters
@@
$N_{\mathrm{max\_val}}$ Maximum number of GPS points in the extremity annulus used for each intersection validation. .
@@
$D_{\mathrm{passing}}$ Maximum distance between the trip and the intersection center for including GPS points in intersection validation.

### v4

@@ Abstract
We propose a new method for inferring roads from GPS trajectories to map construction sites. This task presents a unique challenge due to the erratic and non-standard movement patterns of construction machinery, which significantly diverge from typical vehicular traffic on established roads. Our proposed method first identifies intersections in the road network that serve as critical decision points, and then connects them with edges to produce a graph, which can subsequently be used for planning and task allocation. The method includes several physically meaningful parameters that can be tuned to adapt it to different scenarios. We demonstrate the approach by mapping roads at a real-life construction site in Norway. The method is validated on four increasingly complex segments of the map. In our tests, the method achieved perfect precision and recall in detecting intersections and inferring roads in data with no or low noise, while its performance was reduced in areas with significant noise and consistently missing GPS updates.
@@ 3 Problem definition
The objective is to use GPS data to infer a graph representing the underlying road network, where each node of the graph represents a road intersection or an action point (loading or off-loading) which is described as a vector of latitude, longitude and altitude $I=(lat,lon,altitude)$. An intersection is defined as a point where at least three roads meet. The edges of the graph represent the roads and are defined as a series of consecutive, equidistant points $E=(lat,lon,altitude)$. An example of a graph is shown in Figure 4.
@@ 5 Data preprocessing
    The Haversine formula is applied to convert latitude and longitude into coordinates expressed in meters $(x,y)$, with the origin arbitrarily set to the minimum values of latitude and longitude in the data set (C).
@@
    Each trip is interpolated with a predefined resolution $\Delta D_{\mathrm{interp}}$.
The hyperparameters and their default values are given in E.
Figure 1: Identification of candidate intersections. Upper plot: The GPS trajectories are converted into movement directions $\Phi\in[0,\pi]$ (bi-directional movement, subsection 6.1). Bottom plot: The area is split into grid cells, with the median direction determined for each cell. The median direction is then used to compute the directional dissimilarity $\Delta\phi_{i,j}$ between adjacent cells. The cells with $\Delta\phi_{i,j}\geq\Delta\phi_{\text{thr}}$ are identified as intersection candidates.
@@ 6 Graph inference algorithm
In the first step, the map is partitioned into a grid (Figure 1). Each cell is of size $N_{\mathrm{res}}\times N_{\mathrm{res}}=5\times 5$ m2, and is denoted as $c_{i,j}$, where $i$ and $j$ are the row and column indices, respectively. The GPS trajectories, divided into trips, can then be converted into movement directions $\Phi=\{\phi_{1},\phi_{2},...,\phi_{N}\}$ based on the latitude and longitude of consecutive GPS updates (similar to ijgi4042446). These directions are normalized to a $[0,\pi]$ range (00 to $180^{\circ}$), accounting for bi-directional vehicle movement and treating travel in either direction along the same path as equivalent, i.e.:
$\phi_{\text{norm}}=\begin{cases}\phi&\text{if }0\leq\phi\leq\pi,\\ 2\pi-\phi&\text{if }\pi<\phi<2\pi.\end{cases}$ (1)
Finally, for each set $\Phi_{i,j}=\{\phi_{\text{norm},i,j,1},\phi_{\text{norm},i,j,2},\dots,\phi_{\text{norm},i,j,N}\}$ of points recorded within each cell $c_{i,j}$, the median $\tilde{\phi}_{i,j}$ is computed. This procedure essentially transforms the grid into a 2D histogram, where each cell represents the median of bi-directional vehicular movement.
@@ 6.3 Validation of intersections
The next step consists of filtering out false positives among the candidate intersections using an adapted version of the method proposed in 10.1145/3234692. This method discerns intersections from e.g., road bends by determining the number of roads leaving the candidate location for the intersection. In our version, we first subset the GPS points to those that lie within the circle bounded by the outer radius $R+L$ and apply DBSCAN clustering to those points, with the clustering radius of $\epsilon_{\mathrm{passing}}=12$ m and the minimum sample size $N_{\mathrm{passing}}=5$. The points that are not directly connected to the intersection, i.e., they do not form a cluster with the points passing next to the candidate intersection within a distance of $D_{\mathrm{passing}}=15$ m, are filtered out. Next, following 10.1145/3234692, we define an extremity annulus bounded by two concentric circles with radii $R$ and $R+L$ around each candidate intersection (Figure 2). The GPS points located within the annulus and are connected to the intersection are clustered with a clustering distance of $D_{\mathrm{ext\_clust}}=20$ m. Each cluster of a size larger than $N_{\mathrm{ext\_clust}}=5$ represents a road that passes through the extremity. The final set of intersections are those that have at least three such clusters in their extremity, representing three roads leaving the intersection. The values of $R$ and $L$ can vary depending on the actual size of intersections, and this step can be repeated with multiple values, e.g., $R=[30,100]$ and $L=R+25$.
To reduce computations for determining distances between the GPS points and intersections, we employed k-dimensional trees (k-d trees) that partition space into a hierarchical tree structure, significantly accelerating spatial queries. This approach allowed us to narrow down calculations to only those pairs of GPS points and intersection centers that are within a predefined maximum distance $R_{\mathrm{max}}+L$ from each other. Moreover, while our application did not show significant differences across clustering techniques, HDBSCAN may be a suitable alternative in cases where handling varying densities of GPS data is necessary.
@@ 6.5 Road inference
In the final step of constructing the road network, the nodes, consisting of intersections, loading and drop-off locations, are interconnected by edges that represent the roads. Additional edges are also created to depict roads with a dead end, i.e., roads that pass through only one intersection.
@@
Figure 4: Examples of the inferred graph with crosses representing three types of nodes: intersections (red), load points (green), and drop-off points (black). The actual intersections are represented with orange points. The edges of the graph, i.e., the roads are in blue, while the actual GPS data are represented by gray lines.
Figure 5: Precision and recall for detecting intersections in the considered zones at varying distance thresholds. The detection tolerance radius is the radius around a labeled intersection within which a predicted intersection is considered detected.
7 Results
The visual representation of the performance of the algorithm is presented in Figure 4. Each of the four subplots shows segments of the inferred graph in the order from relatively straightforward to more complex scenarios. The parameters used are defined in E. The intersection validation step was run twice, with $R=30$ m and $R=100$ m (Figure 3).
In the first two examples: Zone 1 and 2, the algorithm achieves perfect recall and precision: All intersections are detected correctly and all driven roads are outlined. The use of at least two radii in the intersection validation step is crucial to detect cross-roads of varying shape and spread, e.g., merging roads (26, Zone 1) and small detours (14, Zone 2). We note that the algorithm is robust against noisy GPS updates present in Zone 1 (visible as straight lines). This robustness against noise that visually resembles a road split is gained when quantifying the directional dissimilarity (subsection 6.2), as the area is never qualified as an intersection candidate (see upper left in Figure 3).
In our further analysis, we show the limitation of the algorithm in the presence of consistent noise. In Zone 3 four intersections are detected correctly (7, 8, 9, and 24). Number 25 is falsely identified as an intersection due to the noisy data. In this case, noisy updates, visible as straight gray lines, are more frequent than the correct outline of the road. Zone 4 presents the biggest challenge for the algorithm, with noisy data dominating several parts of the map, in the areas where vehicles often change direction. In this case, the algorithm correctly detects intersection 6, identifies some intersections near the actual ones (e.g., 21 and 23), and produces several false positives (e.g., 20).
The performance is further quantified with precision and recall (definition in B), as shown in Figure 5. Both metrics improve as the distance threshold increases, reflecting a more relaxed detection criterion. However, the level of alignment between the actual and predicted intersections varies between zones. In Zone 1, high precision and recall ($75\%-100\%$) are achieved even at a threshold of 10–20 m, indicating tight detection. In contrast, Zone 4 requires a threshold of up to 50 m to reach similar values, suggesting a lower level of alignment. Zones 2 and 3 exhibit intermediate behavior. In Zone 2, precision and recall reach $90\%$ at 30 m and $100\%$ at 50 m. For Zone 3, both recall and precision start at $50\%$, eventually reaching $100\%$ recall and $83\%$ precision at at 50 m.
@@ 8 Discussion
    Contextual parameter tuning: The algorithm’s parameters have to be tuned for the specific application and characteristics of the landscape. For instance, the distance parameters that define the relevant area surrounding a singular point marked as an intersection will vary depending on the area covered by an intersection: a well-defined crossroad will need tighter parameter settings than more expansive roundabouts or open-area nodes with several diverging routes.r
@@
    Quality of the GPS data: While multiple problems with the data can be addressed with the steps outlined in section 5, mitigating certain types of noise is not trivial, especially if the inaccurate GPS updates are dominant and overshadow valid data. In such cases, customizing the data cleansing process to the specific deployment context becomes crucial. Additionally, employing techniques to enhance GPS position accuracy can further improve data quality and mitigate the impact of noise aggarwal2020enhancement; kumar2014improving; kumar2013global; kumar2014identifying.
@@ Appendix A Nomenclature
Appendix B Metrics
B.1 Precision and recall
$\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$ (3)
$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (4)
where:
  • TP (True Positives): Predicted intersections that correctly match actual intersections within a defined distance threshold.
  • FP (False Positives): Predicted intersections that do not correspond to any actual intersection.
  • FN (False Negatives): Actual intersections that were not detected by the algorithm.
Appendix C Mathematical formulas
C.1 Haversine formula
The Haversine formula, used for calculating the distance between two points on the surface of a sphere given their latitude and longitude, is given by the following:
$a=\sin^{2}\left(\frac{\Delta\phi}{2}\right)+\cos(\phi_{1})\cos(\phi_{2})\sin^{2}\left(\frac{\Delta\lambda}{2}\right)$
$c=2\arctan 2\left(\sqrt{a},\sqrt{1-a}\right)$
$d=R\cdot c$
where:
  • $\Delta\phi=\phi_{2}-\phi_{1}\quad(\text{latitude difference in radians})$
  • $\Delta\lambda=\lambda_{2}-\lambda_{1}\quad(\text{longitude difference in radians})$
[…]
Appendix D Method flowchart
Preprocessing Convert coordinates into metres. Remove noisy GPS updates. Interpolate paths. Parameters: $D_{\text{endpoints}},\Delta D_{\text{interp}},N_{\text{min}}$ Section 5 2D histograms
Partition the area into a grid. Compute the median heading directions in each cell based on the GPS updates. Parameters: $N_{\text{res}}$
Section 6.1 Intersection identification
Compute the directional dissimilarity in each cell. Flag intersection candidates with the largest dissimilarity. Parameters: $D_{\text{nbr}},D_{\text{int\_clust}},\Delta\phi_{\text{thr}}$
Section 6.2 Intersection validation
Cluster points in extremities. Validate outgoing roads using those clusters. Parameters: $R,L,$ $\epsilon_{\mathrm{passing}}$, $N_{\mathrm{min\_passing}}$, $D_{\text{passing}},D_{\text{ext\_clust}},N_{\text{ext\_clust}}$
Section 6.3 Load/drop-off detection
Cluster load and drop-off locations. Merge points that are close together. Parameters: $D_{\text{load/dump}}$
Section 6.4 Road inference
Connect the nodes with GPS paths. Cluster GPS segments. Assign edges to the nodes. Parameters: $D_{\text{node}},\epsilon_{\text{road}},N_{\text{min\_road}}$
Section 6.5 Final output
Nodes: Intersections, load/drop-off points Edges: Roads connecting nodes
Figure 6: The flowchart of the full road graph generation framework.
Appendix E Hyperparameters for algorithm and data processing
@@
$N_{\mathrm{max\_val}}$ Maximum number of GPS points in the extremity annulus used for each intersection validation.
@@
$\epsilon_{\mathrm{passing}}$ Radius for DBSCAN clustering when validating points within the R+L circles as roads passing through the intersection.
Default: 12 m.
Tuning: Increase to merge more spread-out GPS points. Decrease to avoid merging neighboring but separate roads.
$N_{\mathrm{min\_passing}}$ Minimum sample count for DBSCAN to form a road passing by the intersection (applied to the points within the R+L circles).
Default: 5 points.
$D_{\mathrm{passing}}$ Maximum distance between at least one of the points in each cluster within R+L and the intersection center to include this cluster of points in intersection validation.